How many milliliters of a 5% alcohol solution must be added to 6 milliliters of a 25% alcholol solution to produce a 13% alcohol solution?
Answer by mananth(16946)

percent ---------------- quantity
Alcohol I 25 ---------------- 6 ml
Alcohol II 5 ---------------- x ml
Mixture 13 ---------------- 6 + x ml

25 *6+5x = 13 ( 6 + x )
150 + 5 x = 78 + 13 x
5 x + -13 x = 78 - -150
-8 x = -72
/ -8
x = 9
9 ml of 5 % Alcohol II